# Multi-GraspLLM: A Multimodal LLM for Multi-Hand Semantic Guided Grasp Generation

### v1

@@ Abstract
Multi-hand semantic grasp generation aims to generate feasible and semantically appropriate grasp poses for different robotic hands based on natural language instructions. Although the task is highly valuable, due to the lack of multi-hand grasp datasets with fine-grained contact description between robotic hands and objects, it is still a long-standing difficult task. In this paper, we present Multi-GraspSet, the first large-scale multi-hand grasp dataset with automatically contact annotations. Based on Multi-GraspSet, we propose Multi-GraspLLM, a unified language-guided grasp generation framework. It leverages large language models (LLM) to handle variable-length sequences, generating grasp poses for diverse robotic hands in a single unified architecture. Multi-GraspLLM first aligns the encoded point cloud features and text features into a unified semantic space. It then generates grasp bin tokens which are subsequently converted into grasp pose for each robotic hand via hand-aware linear mapping. The experimental results demonstrate that our approach significantly outperforms existing methods on Multi-GraspSet. More information can be found on our project page https://multi-graspllm.github.io.
@@ 1 Introduction
In this work, we focus on a challenging problem of multi-hand semantic-guided grasp generation. This task aims to generate feasible and semantically appropriate grasp poses for multiple robotic hands through natural language instructions. Unlike traditional approaches that require specific models for each individual robotic hand (Figure 1), this task aims to generate grasp poses of multiple hands with a single model. This unified approach not only provides higher flexibility with only a single model to optimize but also enhances generalization through cross-hand learning. Moreover, while existing methods lack semantic understanding in grasp generation, this task incorporates semantic guidance to generate semantically appropriate grasps based on textual descriptions.
@@
Generally, achieving this goal involves two major challenges. The first challenge is constructing a comprehensive dataset that covers diverse robotic hands and semantic descriptions, which is a non-trivial task. Among existing multi-hand datasets [42, 16, 38, 25] which primarily focus on physically stable grasps, only one dataset [44] provides rich contact semantic information. However, this dataset is limited to the Shadow Hand [31], and its retargeting approach based on human grasp data [50] cannot generalize well to other robotic hands with different link and joint structures, particularly for larger four-fingered systems like Allegro Hand [45]. Moreover, training a unified model for multiple robotic hands faces significant challenges when incorporating semantic guidance, as the same semantic instruction needs to be translated into different grasping behaviors for hands with distinctly different structures. Most semantic-guided methods [44, 43, 15] train a separate model for each hand. While DexGrasp-Diffusion [55] based on DexGraspNet [42] trains a unified diffusion model for different robotic hands, it lacks semantic guidance to grasp objects according to human common sense.
@@
To address the first challenge, we propose Multi-GraspSet, the first large-scale multi-hand grasp dataset with contact annotations. Starting with pre-segmented objects [50, 2], we first generated a large number of physically stable grasps using the unified grasp generation approach [42, 34]. For each grasp case, we then added detailed contact annotations based on the segment annotation. Additionally, we leveraged LLM [1] to generate nearly 1M dialogue samples to support subsequent model training.
@@
For the second challenge, based on the Multi-GraspSet, we design Multi-GraspLLM to generate precise, feasible grasp poses for multiple robotic hands. Multi-GraspLLM leverages the generalization capability and flexibility of large language models to handle variable-length inputs and ouputs, enabling a single model to produce poses for different robotic hands. The model takes a pair of point cloud and text instruction as input and outputs grasp angles for different robotic hands via a hand-aware linear mapping. For example, when given an instruction like “How can we use the Allegro Hand to grasp a glass by its rim?”, Multi-GraspLLM can infer the optimal grasping angles for the Allegro Hand [45] and adapt to other hands trained in the model, such as the Shadow Hand [31] and Panda Gripper [6].
@@
Through extensive experimentation and evaluation of our dataset, we demonstrate that Multi-GraspLLM can successfully generate accurate grasp poses across multiple robotic hand types while maintaining precise control over individual finger positions. Quantitative results show significant improvements in both pose accuracy and grasp quality compared to existing methods. The model exhibits robust generalization capabilities across different grasp configurations while maintaining semantic consistency with natural language instructions.
@@
    We build Multi-GraspSet, the first large-scale multi-hand grasp dataset with contact annotations. Our dataset fills a gap in robotic hand grasp pose generation, particularly in multi-hand and semantic-guided grasping.
@@
    We propose Multi-GraspLLM, an LLM-based multi-hand grasp generation method that integrates LLM with cross-hand robotic grasp generation while ensuring semantically proper grasps.
@@ 2.1 Robotics Grasp Datasets
Grasping datasets can generally be categorized into two types: physically stable and contact-aware. Most existing datasets focus on the former type. DexGraspNet [42, 54] and GenDexGrasp [16] use force-closure [20] as a metric for simulator optimization. Acronym [5] and MultiGripperGrasp [25] using the physical simulator [22] to filter out unstable grasps also support multiple robotic hands but have fewer objects. In contrast, contact-aware datasets transfer human hand poses to robotic hands. DexGYSNet [44] and Human-like-grasp [43] retarget the hand pose from the OakInk [50] dataset to the dexterous robotic hand and its own collected data with segmentation annotations. However, due to the difference in joints and links, DexGYSNet and Human-like-grasp typically only support those robotic hands resembling human hands.
@@ 2.2 Language-guided Grasp Generation
Although extensive research has been conducted on physically-stable grasp generation methods [42, 54, 38, 39, 11, 47, 20, 34, 24], there remains a significant gap in studies focused on language-guided functional grasp generation. The aim is to generate an appropriate grasp pose on a specific functional part of an object based on the description of the grasping scenario. Lerf-togo [30] and graspsplats [9] applied NERF [23] and 3DGS [13] to model 3D semantics scenes, allowing the gripper to grasp the specific object part based on language queries. Human-like-grasp [43] trained a CVAE [33] model that allowed the generation of different functional grasping. However, it cannot support an open vocabulary. GaYs [44] addresses this limitation by training a language-conditioned diffusion model to generate dexterous grasp poses based on language descriptions. At the same time, large models greatly contribute to language-conditioned grasp generation. RealDex [21] leverages Gemini [36] to score generated grasps, filtering out those that satisfy the language description. SemGrasp [15] further extended this idea. It fine-tuned a large language model and used VQ-VAE [40] to discretize low-dimensional hand pose parameters. However, all of these methods only support single-type hand.
@@ 2.3 Multimodal LLMs in Robotics
Multimodal LLMs play a crucial role in reasoning for embodied intelligence, especially with the full development of models such as LLaVA [19] and LLaMA [37]. Several models [29, 46, 12, 48] leverage the reasoning capabilities of vision-language models [19] to identify grasping regions in scene images. Vision-language-action model [14, 41, 57, 17, 10, 18], built on the LLM architecture, enables the generation of discrete actions based on both visual and language input from a scene. Additionally, 3D large language models [51, 32, 35, 49] extend the structure of LLaVA by exchanging the 2D visual encoders with various 3D point cloud encoders such as PointNet [27, 28] or PointBert [52], which enable effective extraction of geometric features from unordered point sets and scene understanding. Methods like [7, 56, 53, 3] refine tasks in embodied environments, while SemGrasp [15] focuses on more specific tasks, displaying human hand grasp poses based on inputs from object point cloud.
@@ 3 Multi-GraspSet Dataset
To our best knowledge, we proposed the first large-scale multi-hand grasp dataset with semantic contact annotation, namely Multi-GraspSet. This dataset was created through unified grasp generation and LLM assistance system. We detail the basic statistics in Sec. 3.1 then introduce the dataset construction process in Sec. 3.2.
@@ 3.1 Dataset statistics
Multi-GraspSet 3 2100 120k 1M
Table 1: Comparison of the contact-aware grasp dataset. Grasp and Con. stand for the number of grasp pose and conversation.
Multi-GraspSet is a large-scale dataset that contains the grasp of three popular robotic hands, 2.1k object point clouds, 120k grasp pose pairs, and more than 1M conversations (Table 1). The dataset features two main types of annotations: contact information and basic conversation. For contact information, we provide precise object contact annotations for each finger, and when all contact-involved fingers interact with the same object part, we provide a more general annotation emphasizing the number of fingers participating in the grasp, as illustrated in Figure 3. Moreover, to enable Multi-GraspLLM to generate grasps following natural language instructions, we construct basic conversations for each grasp pose with different control levels, comprising 5-10 diverse question-answer hand-aware basic conversations focusing on the grasp pose (Figure 6).
Figure 3: Illustration of Multi-GraspSet. Our dataset includes grasp poses of common objects across three robotic hands with two types of contact annotations.
@@ 3.2.1 Unified Grasp Generation
We generate grasp poses of three robotic hand through the unified grasp generation methods (Figure 2). We first collect 2,100 meshes from OakInk [50] and ShapeNet [2], applying convex decomposition to ensure that the meshes are watertight. Since we aim at multi-hand grasp generation, we then use DexGraspNet [42] to generate grasp poses for dexterous robotic hands (Shadow Hand [31] and Allegro Hand [45]), and Contact-GraspNet [34] for grippers (Panda Gripper [6]). DexGraspNet [42] proposes a force-closure method for dexterous hands, working without learning. Contact-GraspNet [34] introduces a new end-to-end grasp generation network for grippers with strong generalization.
@@ 3.2.2 Annotation for Contact Information
where $C(link_{i})$ indicates whether the $i$-th link is in contact with the object, $O$ represents the object’s mesh, and $P$ denotes the point sampled from the robotic hand. Additionally, based on the category labels of object faces, we annotate which specific part of the object the link is in contact with. As briefly mentioned in Sec. 3.1, we use two types of annotation to describe the finger contacts: a detailed annotation that specifies individual fingers - ”Index, middle, thumb, and ring fingers contact the hammer’s grip.”, and a general type for cases if all fingers contact the same part - ”Four fingers grasp the grip of the hammer.”. The general annotation is used to improve text quality and readability when fingers share the same contact location on the object.
@@ 3.2.3 Annotation for Basic Conversation
We categorize basic conversation into three types: low-level, middle-level, and high-level instructions, with levels of detail ranging from coarse to fine. Low-level instructions do not provide contact information and include only basic questions, such as “How do you grasp the {object} using the {hand type}?” Middle-level instructions specify the part to be grasped, for example, “How do you grasp the {part} of the {object} using the {hand type}?”. High-level instructions include all relevant information, such as “Demonstrate the ideal pose of the {hand type} to grasp the {object}: {contact info}.” However, due to the complexity of our contact information compared to CapGrasp [15], fixed templates often lead to awkward phrasing. To address this, we use GPT-4o [1] to refine each sentence, ensuring varied and natural language throughout. Additional details about our annotation system are provided in the Supplementary Material.
@@ 4.1 Multi-GraspLLM
$B=\frac{p-L_{hand}}{W_{hand}},hand\in\{Allegro,Shadow,Panda\}$ (2)
@@
Grasp Pose Generation. Multi-GraspLLM comprises three components: (1) a 3D feature encoder $f_{pc}$ that extract objects geometry information; (2) a modality adaptor $f_{m}$ that aligns object geometry with language features; (3) a large language model $m_{llm}$ for interpreting natural language grasp instructions and generating the grasp pose based on the 3D geometry of object.
Given an object point cloud $P\in\mathbb{R}^{8192}$ and a language description $T$, the point cloud encoder first encodes $P$ into a sequence of feature tokens $s=f_{pc}(P)$, where $s\in\mathbb{R}^{512\times 384}$. The modality adaptor then aligns these point cloud features into the language space $T_{p}=f_{m}(s)$, where $T_{p}\in\mathbb{R}^{512\times 4096}$, to match the language features $T_{l}$ encoded by the LLM language encoder. These two token sequences are concatenated to form a mixed token sequence $T_{m}=[T_{p},T_{l}]$, which serves as input to the LLM backbone. The LLM backbone processes this sequence auto-regressively, generating outputs $O_{i}=m_{llm}(O_{1},...,O_{i-1})$ for the sequence $[O_{1},O_{2},...,O_{n}]$. Finally, the de-tokenizer converts the output sequence into a grasp bin $B=[B_{1},B_{2},...]$. Once the actions are processed into a sequence of bind, Multi-GraspLLM is trained with a standard next-token prediction objective, evaluating the Cross Entropy loss on the predicted grasp bin tokens and generated natural language to enhance the model’s unified generalization capability.
@@
where $[T;R]\in\mathbb{R}^{6}$ represents the 6-D pose of the robotic hand’s wrist, and $\theta\in\mathbb{R}^{d}$ denotes the joint angles specific to different robotic hands ($d=6$ for Panda Gripper [6], $d=22$ for Allegro Hand [45], and $d=28$ for Shadow Hand [31]). The detailed implementation is provided in the supplementary material.

### v3

@@ Abstract
Multi-hand semantic grasp generation aims to generate feasible and semantically appropriate grasp poses for different robotic hands based on natural language instructions. Although the task is highly valuable, due to the lack of multi-hand grasp datasets with fine-grained contact description between robotic hands and objects, it is still a long-standing difficult task. In this paper, we present Multi-GraspSet, the first large-scale multi-hand grasp dataset with automatically contact annotations. Based on Multi-GraspSet, we propose Multi-GraspLLM, a unified language-guided grasp generation framework, which leverages large language models (LLM) to handle variable-length sequences, generating grasp poses for diverse robotic hands in a single unified architecture. Multi-GraspLLM first aligns the encoded point cloud features and text features into a unified semantic space. It then generates grasp bin tokens that are subsequently converted into grasp pose for each robotic hand via hand-aware linear mapping. The experimental results demonstrate that our approach significantly outperforms existing methods in both real-world experiments and simulator. More information can be found on our project page https://multi-graspllm.github.io.
@@ 1 Introduction
In this work, we focus on a challenging problem of multi-hand semantic-guided grasp generation. This task aims to generate feasible and semantically appropriate grasp poses for multiple robotic hands through natural language instructions. Unlike traditional approaches that require specific models for each individual robotic hand (Figure 1), this task aims to generate grasp poses of multiple hands with a single model. The unified approach not only offers greater flexibility by optimizing a single model when deploying, but also enhances generalization of grasp generation models through cross-hand learning. Furthermore, while existing methods lack semantic understanding in grasp generation, this task incorporates semantic guidance to generate semantically appropriate grasps based on textual descriptions.
@@
Generally, achieving this goal involves two major challenges. The first challenge is to build a comprehensive dataset that includes various robotic hands and semantic descriptions. However, this is not a trivial task. Among existing multi-hand datasets [49, 19, 45, 30] that focus mainly on physically stable grasps, only one dataset [52] provides rich contact semantic information. However, its retargeting approach based on human grasp data [58] does not generalize well to robotic hands with different link and joint structures, especially multi-fingered systems with fewer than five fingers, such as the Jaco Hand [11] and Barrett Hand [2]. Moreover, training a unified model for multiple robotic hands faces additional challenges when incorporating semantic guidance, as the same semantic instruction must be translated into different grasping behaviors for hands with distinctly different robotic structures. Most semantic-guided methods [52, 50, 18] train a separate model for each hand. Although DexGrasp-Diffusion [63] based on DexGraspNet [49] trains a unified diffusion model for different robotic hands, it lacks semantic guidance to grasp objects.
@@
To address the first challenge, we propose Multi-GraspSet, the first large-scale multi-hand grasp dataset with contact annotations. Starting with pre-segmented objects [58, 3], we first generated a large number of physically stable grasps using the unified grasp generation approach [49, 41, 28]. For each grasp case, we then added detailed contact annotations based on the segment annotation. Furthermore, we used LLM [1] to generate nearly 1.2M dialogue samples to support subsequent model training.
@@
For the second challenge, based on the Multi-GraspSet, we design Multi-GraspLLM to generate precise, feasible grasp poses for multiple robotic hands. Multi-GraspLLM leverages the generalization capability and flexibility of large language models to handle variable-length inputs and outputs, enabling a single model to produce poses for different robotic hands. The model takes a pair of point cloud and text instruction as input and outputs grasp angles for different robotic hands via a hand-aware linear mapping. For example, given the instruction “How can we use the Allegro Hand to grasp a glass by its rim?”, Multi-GraspLLM can infer the optimal grasp angles for the Allegro Hand [53]. Similarly, when asked “How can we use the Shadow Hand to grasp a glass by its rim?”, the same model can infer the appropriate grasp angles for the Shadow Hand [38].
@@
Through extensive experiments and evaluation of our dataset, we demonstrate that Multi-GraspLLM can successfully generate accurate grasp poses across multiple types of robotic hands while maintaining precise control over individual finger positions. Quantitative results in simulator and real-world experiments show significant improvements in both pose accuracy and grasp quality compared to existing methods. The model exhibits robust generalization capabilities across different grasp configurations while maintaining semantic consistency with natural language instructions.
@@
    We build Multi-GraspSet, the first large-scale multi-hand grasp dataset with rich contact annotations. Our dataset could help fill a gap in robotic hand grasp pose generation, particularly in multi-hand and semantic-guided grasping.
@@
    We propose Multi-GraspLLM, an LLM-based multi-hand semantic grasp generation method that enables cross-hand robotic grasp generation while ensuring semantically proper grasps. Our work could also inspire related research of embodied foundation models for different robotics systems.
  3. 3.
    Experimental results show that Multi-GraspLLM achieves state-of-the-art performance in grasp generation. Moreover, joint training of a multi-hand model yields better results than training separate models for each hand.
@@ 2.1 Robotics Grasp Datasets
Grasping datasets can generally be categorized into two types: physically stable and contact-aware. Most existing datasets focus on the former type. DexGraspNet [49, 62] and GenDexGrasp [19] use force-closure [23] as a metric for simulator optimization. Acronym [6] and MultiGripperGrasp [30] using the physical simulator [25] to filter out unstable grasps also support multiple robotic hands but have fewer objects. In contrast, contact-aware datasets transfer human hand poses to robotic hands. DexGYSNet [52] and Human-like-grasp [50] retarget the hand pose from the OakInk [58] dataset to the dexterous robotic hand and its own collected data with segmentation annotations. However, due to the difference in joints and links, DexGYSNet [52] and Human-like-grasp [50] typically only support robotic hands resembling human hands.
@@ 2.2 Language-guided Grasp Generation
Although extensive research has been conducted on physically-stable grasp generation methods [49, 62, 45, 46, 14, 55, 23, 41, 29], there remains a significant gap in studies focused on language-guided functional grasp generation. The aim is to generate an appropriate grasp pose on a specific functional part of an object based on the description of the grasping scenario. LERF-TOGO [35] and GraspSplats [12] applied NERF [27] and 3DGS [16] to model 3D semantics scenes, allowing the gripper to grasp the specific object part based on language queries. Human-like-grasp [50] trained a CVAE [40] model that allowed the generation of different functional grasping. However, it cannot support an open vocabulary. DexGYS [52] addresses this limitation by training a language-conditioned diffusion model to generate dexterous grasp poses based on language descriptions. At the same time, large models greatly contribute to language-conditioned grasp generation. RealDex [24] leverages Gemini [43] to score the generated grasps, filtering out those that satisfy the language description. SemGrasp [18] further extended this idea. It fine-tuned a large language model and used VQ-VAE [47] to discretize low-dimensional hand pose parameters. However, all of these methods only support a single type of hand.
@@ 2.3 Multimodal LLMs in Robotics
Multimodal LLMs play a crucial role in reasoning for embodied intelligence, especially with the full development of models such as LLaVA [22] and LLaMA [44]. Several models [34, 54, 15, 56] leverage the reasoning capabilities of vision-language models [22] to identify grasping regions in scene images. Vision-language-action model [17, 48, 65, 20, 13, 21, 26], built on the LLM architecture, enables the generation of discrete actions based on both visual and language input from a scene. Additionally, 3D large language models [59, 39, 42, 57] extend the structure of LLaVA by exchanging the 2D visual encoders with various 3D point cloud encoders such as PointNet [32, 33] or PointBERT [60], which enable effective extraction of geometric features from unordered point sets and scene understanding. Methods like [8, 64, 61, 4] refine tasks in embodied environments, while SemGrasp [18] focuses on more specific tasks, displaying human hand grasp poses based on inputs from object point cloud.
@@ 3 Multi-GraspSet Dataset
To our best knowledge, we proposed the first large-scale multi-hand grasp dataset with semantic contact annotation, namely Multi-GraspSet. This dataset was created through unified grasp generation and LLM assistance system. We detail the basic statistics in Sec. 3.1, and then introduce the dataset construction process in Sec. 3.2.
@@ 3.1 Dataset Statistics
Multi-GraspSet 5 2100 140k 1.1M
Table 1: Comparison of the contact-aware grasp dataset. “Grasp” and “Con.” stand for the number of grasp pose and conversation.
Multi-GraspSet is a large-scale dataset that contains the grasp of five popular robotic hands, 2.1k object point clouds, 140k grasp pose pairs, and more than 1M conversations (Table 1). The dataset features two main types of annotations: contact information and basic conversation. For contact information, we provide precise object contact annotations for each finger, and when all contact-involved fingers interact with the same object part, we provide a more general annotation emphasizing the number of fingers participating in the grasp, as illustrated in Figure 3. Moreover, to enable Multi-GraspLLM to generate grasps following natural language instructions, we construct basic conversations for each grasp pose with different control levels, comprising 5-10 diverse question-answer hand-aware basic conversations focusing on the grasp pose (Figure 6).
Figure 3: Illustration of Multi-GraspSet. Our dataset includes grasp poses of common objects across five robotic hands with two types of contact annotations.
@@ 3.2.1 Unified Grasp Generation
We generate grasp poses of five robotic hands through the unified grasp generation methods [49, 28, 41] (Figure 2). Specifically, DexGraspNet [49] and GraspIt [28] proposes an analysis-based method for dexterous hands, working without learning. Contact-GraspNet [41] introduces an end-to-end grasp generation network for grippers with strong generalization. We first collect 2,100 meshes from OakInk [58] and ShapeNet [3], applying convex decomposition to ensure that the meshes are watertight. Since we aim at multi-hand grasp generation, we then use DexGraspNet [49] and GraspIt [28] to generate grasp poses for dexterous robotic hands (Shadow Hand [38], Allegro Hand [53], Barrett Hand [2] and Jaco hand [11]), and Contact-GraspNet [41] for grippers (Panda Gripper [7]).
@@ 3.2.2 Annotation for Contact Information
where $C(link_{i})$ indicates whether the $i$-th link is in contact with the object, $O$ represents the object’s mesh, and $P$ denotes the point sampled from the robotic hand. Additionally, based on the category labels of object faces, we annotate which specific part of the object the link is in contact with. As briefly mentioned in Sec. 3.1, we use two types of annotation to describe the finger contacts: a detailed annotation that specifies individual fingers—“Index, middle, thumb, and ring fingers contact the hammer’s grip” and a general type for cases where all fingers contact the same part—“Four fingers grasp the grip of the hammer.” The general annotation is used to improve text quality and readability when fingers share the same contact location on the object.
@@ 3.2.3 Annotation for Basic Conversation
We categorize basic conversation into three types: low-level, middle-level, and high-level instructions, with levels of detail ranging from coarse to fine. Low-level instructions do not provide contact information and include only basic questions (e.g., “How do you grasp the {object} using the {hand type}?”) Middle-level instructions specify the part to be grasped (e.g., “How do you grasp the {part} of the {object} using the {hand type}?”) High-level instructions include all relevant information (e.g., “Demonstrate the ideal pose of the {hand type} to grasp the {object}: {contact info}.”) However, due to the complexity of our contact information compared to CapGrasp [18], fixed templates often lead to awkward phrasing. To address this, we use GPT-4o [1] to refine each sentence, ensuring varied and natural language throughout. More details about our annotation system can be found in the Supplementary Material.
@@ 4.1 Multi-GraspLLM
$B=\frac{p-L_{hand}}{W_{hand}},hand\in\{Allegro,Shadow,\cdots\}$ (2)
@@
Grasp Pose Generation. Multi-GraspLLM comprises three components: 1) a 3D feature encoder $f_{pc}$ that extract objects geometry information; 2) a modality adaptor $f_{m}$ that aligns object geometry with language features; 3) a large language model $m_{llm}$ to interpret natural language grasp instructions and generate the grasp pose based on the 3D geometry of object.
Given an object point cloud $P\in\mathbb{R}^{8192}$ and a language description $T$, the point cloud encoder first encodes $P$ into a sequence of feature tokens $s=f_{pc}(P)$, where $s\in\mathbb{R}^{512\times 384}$. The modality adaptor then aligns these point cloud features into the language space $T_{p}=f_{m}(s)$, where $T_{p}\in\mathbb{R}^{512\times 4096}$, to match the language features $T_{l}$ encoded by the LLM language encoder. These two token sequences are concatenated to form a mixed token sequence $T_{m}=[T_{p},T_{l}]$, which serves as input to the LLM backbone. The LLM backbone processes this sequence auto-regressively, generating outputs $O_{i}=m_{llm}(O_{1},...,O_{i-1},T_{m})$ for the sequence $[O_{1},O_{2},...,O_{n}]$. Finally, the de-tokenizer converts the output sequence into a grasp bin $B=[B_{1},B_{2},...]$. Each bin acts as a special token, analogous to a word in a large language model. Each bin serves as a special token, analogous to the output of words in LLM. Once the actions are processed into a sequence of bins, Multi-GraspLLM is trained with a standard next-token prediction objective, evaluating the Cross Entropy loss on the predicted grasp bin tokens and generated natural language to enhance the model’s unified generalization capability.
@@
where $[T;R]\in\mathbb{R}^{6}$ represents the 6-D pose of the robotic hand’s wrist, and $\theta\in\mathbb{R}^{d}$ denotes the joint angles specific to different robotic hands (e.g , $d=22$ for Allegro Hand [53] as shown in the Fig. 4). The detailed implementation is provided in the supplementary material.